OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Sn:    2 + 5 + 8 + . . . + ( 3n - 1) = n(1 + 3n)/2 I Got S1: 2 = 1(3 ・ 1 + 1) /2 2 =1 ・ 4 / 2 2 = 2 My teacher says my calculations have arithmetic errors

OpenStudy (anonymous):

@ybarrap could you help me out please

OpenStudy (ybarrap):

S1= 1(1 + 3)/2= 4/2 = 2 S2 = 2(1 + 3*2)/2 = 7 S3 = 3(1 + 3*3)/2=3(10)/2 = 15 So, S1=2, S2=7 & S3=15. Does this make sense?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thank You :)

OpenStudy (ybarrap):

NP

OpenStudy (ybarrap):

Hey - Vote for my art if you like it - http://openstudy.com/study#/updates/52e2b272e4b096fa40590b00